OpenStudy (anonymous):

Tom borrowed $35,000 to remodel his house. At the end of the 5-year loan, he had repaid a total of $46,375. At what simple interest rate did he borrow the money?

OpenStudy (anonymous):

total paid = p * i * n p = principal (original amount borrowed or loaned) i = interest rate for one period n = number of periods 35,000 * x * 5 = 46,375 175000x = 46,375 = .265 .265*100 = 26.5%